# Rotationally symmetric translating solitons of fully nonlinear extrinsic geometric flows: Classification and Applications

### v1

@@ 6. Uniqueness of rotationally symmetric graphs and Applications
where $a=\alpha\bigl(\partial_{x}\tilde{\gamma}(0,1)-\beta\bigr)\in\mathbb{R}$ [^6] When $\alpha=1$, $a=\partial_{x}\tilde{\gamma}(0,1)$.. Then, for any $C>0$ there exists $R\gg 1$ such that
@@ 7. Conclusions
These constructions provide a framework for constructing $\Delta$-wing solitons for non-degenerate (n1)(n-1) normalized curvature functions, following the variational approaches employed by Hoffman-Ilmanen-Martin-White [hoffman2019graphical] as well as the more EDP-based techniques of Bourni-Langford-Tinaglia [bourni2020existence] under the condition of being rotationally symmetric at infinity, since such solitons arise naturally from the process of desingularizations between bowl-type translators and grim-reaper cylinders. Furthermore, the catenoidal translators we construct should allow stability results for complete graphical solutions of the associated $\gamma$-flow, in the spirit of Clutterbuck-Schnürer-Schulze [clutterbuck2007stability].

### v2

@@ 6. Uniqueness of rotationally symmetric graphs and Applications
where $a=\alpha\bigl(\partial_{x}\tilde{\gamma}(0,1)-\beta\bigr)$. Then, for any $C>0$ there exists $R\gg 1$ such that
@@ 7. Conclusions
These constructions provide a framework for constructing $\Delta$-wing solitons for non-degenerate (n1)(n-1) normalized curvature functions, following the variational approaches employed by Hoffman-Ilmanen-Martin-White [hoffman2019graphical], as well as the more EDP-based techniques of Bourni-Langford-Tinaglia [bourni2020existence] under the condition of being rotationally symmetric at infinity, since such solitons arise naturally from the desingularization process between the bowl soliton and grim-reaper cylinders for the mean curvature function $\gamma=H$. Furthermore, the catenoidal translators we construct should allow stability results for complete graphical solutions of the associated $\gamma$-flow, in the spirit of Clutterbuck-Schnürer-Schulze [clutterbuck2007stability].